OpenStudy (anonymous):

The ratio of the areas of two similar polygons is 9:4. If the perimeter of the first polygon is 17 cm, what is the perimeter of the second polygon? Round to the nearest tenth.

OpenStudy (anonymous):

Upscale the ratio by a factor of 17/9: 9:4 = (17/9)*9:(17/9)*4 = 17:(17/9)*4 = 7.55555... (recurring) To the nearest tenth, the answer is 7.6

OpenStudy (anonymous):

A. 13cm B. 10.5cm C. 11.3cm D. 15.4